Pls assist in solving:+what+number+must+be+added+to+each+term+of+the++11+ratio+15+so+that+it+becomes+7+ratio+8
Answer by stanbon(75887)

what number must be added to each term of the 11 ratio 15 so that it becomes 7 ratio 8
-----
Equation::
(11+x)/(15+x) = 7/8
------------
88+8x = 135 + 7x
x = 47
------------------
Cheers,
Stan H.
------------------